OpenStudy (i_love_my_nieces):

Help! What is the slope of the line that goes though (6, 5) and (3, 1)? 1. 3/4 2. -3/4 3. 4/3 4. -4/3 @SolomonZelman

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf m=~~ \frac{y_1-y_2}{x_1-x_2} }\) can you plug in your points ?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf m=~~ \frac{y_1-y_2}{x_1-x_2} }\) to plug in (6, 5) and (3, 1) say that ( 6 , 5 ) \(\large\color{blue}{ \bf (x_1~,~y_1) }\) ( 3 , 1 ) \(\large\color{blue}{ \bf (x_2~,~y_2) }\) seee what I mean ?

OpenStudy (solomonzelman):

6=x1 3=x2 5=y1 1=y2

OpenStudy (i_love_my_nieces):

Yeah

OpenStudy (i_love_my_nieces):

@SolomonZelman

OpenStudy (solomonzelman):

Please show me how YOU would plug in the points into the slope formula.

OpenStudy (i_love_my_nieces):

I showed you!!

OpenStudy (solomonzelman):

I don't see any of your work though -:(

OpenStudy (i_love_my_nieces):

Like how you did it up there ^^^^^^

OpenStudy (i_love_my_nieces):

Idk how to work the equation box

OpenStudy (solomonzelman):

You said that you showed me how you plugged in the points into the slope formula... and I don't see any work posted by you. You can draw using the drawing tool, or the equation editor, \frac{ numerator }{ denominator } is a fraction latex x1 and x2 are x_1 x_2 y1 and y2 are y_1 y_2

OpenStudy (i_love_my_nieces):